(4-7i)(4+7i)
please solve and show work thank you

Answer by Alan3354(69443)

(4-7i)(4+7i)
------------
= 4*(4+7i) - 7i*(4+7i)
= 16 + 28i - 28i - 49i^2
= 16 - 49i^2
i^2 = -1
= 16 + 49
= 65
------------
These are similar to (a-b)*(a+b) = a^2 - b^2 EXCEPT the sign is plus.
(a - bi)*(a + bi) = a^2 + b^2
------
Also, nothing is solved for, there's no variable in 65. It's only simplified.